DIPOLE MOMENT AND ROVIBRATIONAL INTENSITIES IN THE ELECTRONIC GROUND STATE OF NH_{3}: BRIDGING THE GAP BETWEEN AB INITIO THEORY AND SPECTROSCOPIC EXPERIMENT.

SERGEI N. YURCHENKO, JINGJING ZHENG, WALTER THIEL, Max-Planck-Institut für Kohlenforschung, Kaiser-Wilhelm-Platz 1, D–45470 Mülheim an der Ruhr, Germany; MIGUEL CARVA-JAL, Departamento de Fisica Aplicada, Facultad de Ciencias Experimentales, Avda. de las FF.AA. s/n, Universidad de Huelva, 21071, Huelva, Spain; HAI LIN, Department of Chemistry, University of Minnesota, 207 Pleasant St. SE. Minneapolis, Minnesota 55455; \underline{PER JENSEN}, FB C – Theoretische Chemie, Bergische Universität, D–42097 Wuppertal, Germany.

We report theoretical values for the transition moments of an extensive set of vibrational bands in the electronic ground state of ^{14}NH_{3}. For selected bands, we have further made detailed simulations of the rotational structure. The calculations are carried out by means of recently developed computational procedures for describing the nuclear motion^{a} and are based on a high-level ab initio potential energy surface, and high-level dipole moment surfaces, for the electronic ground state of NH_{3}. The reported theoretical intensity values are compared to, and found to agree very well with, corresponding experimental results. It is believed that the computational method, in conjunction with high-quality ab initio potential energy and dipole moment surfaces, can simulate rotation-vibration spectra of XY_{3} pyramidal molecules prior to observation with sufficient accuracy to facilitate the observation of t hese spectra. By degrading the accuracy of selected elements of the calculations, we have also investigated the influence of customary approximations on the computed intensity values.

aS. N. Yurchenko, M. Carvajal, P. Jensen, H. Lin, J. Zheng, and W. Thiel, Mol. Phys. 103, 359-378 (2005) and references therein.